Circular No. 8333

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2004 HC_{18} (LINEAR) An apparently asteroidal object discovered by the LINEAR survey

(discovery observation given below from MPS 105288) has been found to have cometary appearance by A. Gleason and J. Larsen on CCD exposures taken by J. Scotti with the Spacewatch telescope on Apr. 25.34–25.39 UT, when it showed a coma of diameter ∼ 5′′ (and mag 17.6–17.9) and two tails: a diffuse component ≈ 7′′ long in p.a. 350^{o}–355^{o} and a very diffuse but more obvious tail ≈ 10′′–16′′ long spanning p.a. 275^{o}–335^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Apr. 17.34932  15^{h}10^{m}51.^{s}35  −19^{o}15′39′′.8  19.1

Additional astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2004-H90.

T = 2004 June 18.494 TT  ω = 30.^{o}716
e = 0.51050  Ω = 219.568 }2000.0
q = 1.71315 AU  i = 23.542
a = 3.49981 AU n^{o} = 0.150535 P = 6.55 years

COMETS C/2003 K13, C/2003 M8–M12, C/2003 N1–N2 (SOHO) Further to IAUC 8327, K. Battams reports his measurements of the

positions of several Kreutz sungrazing comets found on SOHO C2 website images by D. Evans (C/2003 K13), X.-m. Zhou (C/2003 M8, M9, M10), X. Leprette (C/2003 M11, M12, N1), R. Kracht (C/2003 M12), and J. Sachs (C/2003 N1, N2). C/2003 N1 was also observed on C3 images. The reduced astrometry by B. G. Marsden and R. Kisala and orbital elements appear on the MPECs cited below.

Comet  2003 UT  α_{2000}  δ_{2000}  MPEC
C/2003 K13 May 23.910  4^{h}02.^{m}3  +19^{o}01′  2004-H49
C/2003 M8 June 22.246  5 56.8  +21 39  2004-H49
C/2003 M9  23.169  6 00.7  +22 03  2004-H49
C/2003 M10  23.188  6 00.3  +21 48  2004-H49
C/2003 M11  23.329  6 01.4  +21 35  2004-H49
C/2003 M12  23.421  6 01.1  +21 41  2004-H49
C/2003 N1  July 5.763  6 49.8  +21 29  2004-H87
C/2003 N2  13.029  7 16.0  +20 11  2004-H87
2004 April 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green